# A Comprehensive Survey of Small Language Models in the Era of Large Language Models: Techniques, Enhancements, Applications, Collaboration with LLMs, and Trustworthiness

### v1

@@ Abstract.
Large language models (LLM) have demonstrated emergent abilities in text generation, question answering, and reasoning, facilitating various tasks and domains. Despite their proficiency in various tasks, LLMs like LaPM 540B and Llama-3.1 405B face limitations due to large parameter sizes and computational demands, often requiring cloud API use which raises privacy concerns, limits real-time applications on edge devices, and increases fine-tuning costs. Additionally, LLMs often underperform in specialized domains such as healthcare and law due to insufficient domain-specific knowledge, necessitating specialized models. Therefore, Small Language Models (SLMs) are increasingly favored for their low inference latency, cost-effectiveness, efficient development, and easy customization and adaptability. These models are particularly well-suited for resource-limited environments and domain knowledge acquisition, addressing LLMs’ challenges and proving ideal for applications that require localized data handling for privacy, minimal inference latency for efficiency, and domain knowledge acquisition through lightweight fine-tuning. The rising demand for SLMs has spurred extensive research and development. However, a comprehensive survey investigating issues related to the definition, acquisition, application, enhancement, and reliability of SLM remains lacking, prompting us to conduct a detailed survey on these topics. The definition of SLMs varies widely, thus to standardize, we propose defining SLMs by their capability to perform specialized tasks and suitability for resource-constrained settings, setting boundaries based on the minimal size for emergent abilities and the maximum size sustainable under resource constraints. For other aspects, we provide a taxonomy of relevant models/methods and develop general frameworks for each category to enhance and utilize SLMs effectively. We have compiled the collected SLM models and related methods on GitHub: https://github.com/FairyFali/SLMs-Survey.
@@ 1. Introduction
The evolution of neural language models (LMs) from BERT’s (Devlin et al., 2019) pre-training and fine-tuning paradigm to T5’s (Raffel et al., 2020) pre-training plus prompting approach, and finally to GPT-3’s (Brown et al., 2020) pre-training plus in-context learning, has greatly enhanced natural language processing (NLP). These advancements have broadened NLP’s application across various fields, including language understanding (Wang et al., 2018), programming (Nam et al., 2024; Team, 2024a), recommendation systems (Wang et al., 2024f), information retrieval (Spatharioti et al., 2023; Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020), mobile-device control (Ding, 2024), scientific discovery (Shojaee et al., 2024; Zhang et al., 2024e), medical question answering (Wang et al., 2023d; Bolton et al., 2024b), and legal question answering (Almeida et al., 2024). In particular, the recent emergence of proprietary commercial models including ChatGPT, Bard, and Claude, and open-sourced models such as Llama (Touvron et al., 2023a, b; Dubey et al., 2024) has led to rapid growth in the development of large language models (LLMs). Even though neural networks consistently improve on various tasks with longer training times, larger datasets, and increased model sizes—a phenomenon known as a neural scaling law (Kaplan et al., 2020), these models unpredictably exhibit a sudden acquisition of versatile abilities, termed ”emergent ability,” once they reach a critical scale threshold, thereby supporting the ”larger is better” trend. This ability is not present in small-scale models. For instance, the latest Llama-3.1 model with 405 billion parameters performs better in dialogue, logical reasoning, and programming compared to the smaller 7B counterpart (Dubey et al., 2024).
@@
Despite their prowess in complex tasks, LLMs’ huge parameters and computational needs impose significant limitations, hindering their adoption in many real-world applications. For example, the LLaMa 3.1 model with 405 billion parameters (Dubey et al., 2024), trained on 16K H100 GPUs for 54 days, requires about 202.5 GB of GPU memory using int4 precision and has large inference latency. These issues present several challenges in specific contexts: (1) LLMs are generally hosted in the cloud and used via cloud-based APIs due to the large GPU memory and computational cost. Users need to upload their data to query LLMs, raising data leakage and privacy concerns, especially in high-stake scenarios such as healthcare, finance, and e-commerce; (2) Driven by personal agents, on-device deployment is a critical requirement. Several factors, including cloud costs, latency, and privacy concerns, hinder the on-device processing of cloud-based LLMs, and direct deployment is impractical due to their high parameter and cache requirements, which often exceed the capabilities of devices such as mobile phones; (3) Their large parameter count can cause inference delays from seconds to minutes, unsuitable for real-time applications. For instance, Llama 2 7B takes approximately 84 seconds to process 100 tokens on benchmarks including HellaSwag, TruthfulQA, MMLU, and Arc_C when run on a smartphone equipped with a Snapdragon 685 processor (Thawakar et al., 2024); (4) To boost performance in specialized domains like healthcare and law, where generic LLMs underperform, LLMs are often fine-tuned. However, this process is computationally expensive due to their large size. (5) Though general-purpose LLMs are powerful, many real-world applications require only specific abilities and domain knowledge, deploying general-purpose LLMs would be a waste of resources and such LLMs often cannot match the performance of models tailored for specific tasks (Phogat et al., 2024; Guo et al., 2023a; Jeong et al., 2023; Wang et al., 2024f; Chen et al., 2024a).
@@
Recently, small language models (SLMs) have shown great potential in alleviating these issues while achieving performance comparable to LLMs for domain-specific problems (Yang et al., 2024e; Team et al., 2024b; Bellagente et al., 2024; Pfeiffer et al., 2024; Zhang et al., 2024i; Groeneveld et al., 2024; Hu et al., 2024b; Abdin et al., 2024; Thawakar et al., 2024; Liu et al., 2024h). We Owing to fewer parameters, SLMs excel in efficiency, cost, flexibility, and customization. They provide significant computational savings in pre-training and inference with reduced memory and storage needs, which is vital for applications requiring efficient resource use. These small models are especially effective in resource-limited settings, performing well on low-power devices such as edge devices. Besides, SLMs improve on-device processing by enhancing privacy, security, response times, and personalization. This supports advanced personal assistants and cloud-independent applications, boosting energy efficiency and reducing carbon emissions. For example, the Llama 3.2 models (1B & 3B) demonstrate that local processing enables immediate execution of prompts and responses (AI, 2024). This approach protects privacy by keeping sensitive data such as patient health information (PHI), business data, personal messages, and calendar details local, enhancing confidentiality. It also allows for precise control over which queries are processed on-device versus those requiring cloud-based models. Therefore, small language models are gaining increasing attention as alternatives to LLMs, as indicated in Figure 2, which shows that SLMs are downloaded more frequently than larger models in the Hugging Face community, and Figure 3, which illustrates the growing popularity of SLM releases over time.
@@
Typically, LMs that exhibit emergent abilities are classified as LLMs. However, the categorization of SLMs remains unclear. Studies vary in their contexts: some define SLMs as models with fewer than one billion parameters (Liu et al., 2024h), while others consider the term “small language model” relative to the larger counterparts (Wang et al., 2024f; Tang et al., 2024b; Lee et al., 2024), with no consensus on a unified definition in the current landscape of LLMs. Research suggests SLMs for mobile devices, typically possessing around 6GB of memory, consist of sub-billion parameter models (Liu et al., 2024h), whereas others classify models with up to 10 billion parameters as small, noting their lack of emergent abilities (Fu et al., 2023). Given their use in resource-constrained environments and for specific tasks, we propose a generalized definition: Given specific tasks and resource constraints, we define SLMs as falling within a range where the lower bound is the minimum size at which the model exhibits emergent abilities for a specialized task, and the upper bound is the largest size manageable within limited resource conditions. This definition integrates various perspectives and addresses factors related to mobile computing and capability thresholds.
@@
Due to the growing demand for SLMs, extensive literature has emerged on various aspects of SLMs. For example, several training techniques optimized for SLMs, such as quantization-aware training (Xu et al., 2024; Wang et al., 2023e; Liu et al., 2023a) and selective architectural component choices (Thawakar et al., 2024; Liu et al., 2024h), aim to enhance performance in specific applications (Phogat et al., 2024; Roziere et al., 2023; Wu et al., 2024d; Bonifacio et al., 2022; Chen et al., 2024a). These methods have led to the development of numerous open-source, general-purpose, and domain-specific SLMs (Yang et al., 2024e; Team et al., 2024b; Bellagente et al., 2024; Acikgoz et al., 2024; Bolton et al., 2024b; Zhang et al., 2024d). Beyond their inherent capabilities, SLMs can also serve as a module or effective proxies for enhancing LLMs (Zhao et al., 2023b; Xu et al., 2023a; Wu et al., 2024a; Sha and Zhang, 2024; Mitchell et al., 2024; Yang et al., 2024d). Despite the commendable performance of SLMs, it is crucial not to overlook their credibility issues, such as the risks of producing hallucinations and privacy breaches (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024). However, currently, there is no comprehensive survey thoroughly exploring these works on SLMs in the era of LLMs. Therefore, this paper offers the first comprehensive survey that analyzes various aspects of SLMs in the LLM era and their future directions. The overview structure of our paper is shown in Figure 1. To summarize, our major contributions are:
@@
Figure 4. Transformer architecture.

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated emergent abilities in text generation, question answering, and reasoning, facilitating various tasks and domains. Despite their proficiency in various tasks, LLMs like PaLM 540B and Llama-3.1 405B face limitations due to large parameter sizes and computational demands, often requiring cloud API use which raises privacy concerns, limits real-time applications on edge devices, and increases fine-tuning costs. Additionally, LLMs often underperform in specialized domains such as healthcare and law due to insufficient domain-specific knowledge, necessitating specialized models. Therefore, Small Language Models (SLMs) are increasingly favored for their low inference latency, cost-effectiveness, efficient development, and easy customization and adaptability. These models are particularly well-suited for resource-limited environments and domain knowledge acquisition, addressing LLMs’ challenges and proving ideal for applications that require localized data handling for privacy, minimal inference latency for efficiency, and domain knowledge acquisition through lightweight fine-tuning. The rising demand for SLMs has spurred extensive research and development. However, a comprehensive survey investigating issues related to the definition, acquisition, application, enhancement, and reliability of SLM remains lacking, prompting us to conduct a detailed survey on these topics. The definition of SLMs varies widely, thus to standardize, we propose defining SLMs by their capability to perform specialized tasks and suitability for resource-constrained settings, setting boundaries based on the minimal size for emergent abilities and the maximum size sustainable under resource constraints. For other aspects, we provide a taxonomy of relevant models/methods and develop general frameworks for each category to enhance and utilize SLMs effectively. We have compiled the collected SLM models and related methods on GitHub: https://github.com/FairyFali/SLMs-Survey.
@@ 1. Introduction
for tree= draw, shape=rectangle, rounded corners, top color=white, grow’=0, l sep’=1.2em, reversed=true, anchor=west, child anchor=west, , forked edges, root/.style= draw=customred, rotate=90, shading angle=90, bottom color=white!40, anchor=north, font=, inner sep=0.5em, level1/.style= draw=customred, text width=2cm, bottom color=white!30, font=, inner sep=0.3em, s sep=0.2em, level2/.style= draw=customred, text width=4cm, bottom color=white!40, font=, inner sep=0.25em, s sep=0.1em, level3/.style= draw=customred, text width=4cm, bottom color=white!40, font=, inner sep=0.2em, l sep’=0.5em, level4/.style= draw=customred, text width=6.7cm, top color=customyellow!100, bottom color=customyellow!100, font=, inner sep=0.2em, l sep’=0.5em, where n=0root, where level=1level1, where level=2level2, where level=3level3, where level¿=4level4, [Small Language Models [Introduction (§1) ] [Concepts in Building LMs (§2) [Architecture (§2.1)] [Training techniques (§2.2)] [Obtain SLMs from LLMs (§2.3) [Pruning (§2.3.1) [Unstructured Pruning (Frantar and Alistarh, 2023; Sun et al., 2024c; Zhang et al., 2023a, 2024a; Li et al., 2023f; Shao et al., 2024; Das et al., 2023); Structured Pruning (Men et al., 2024; Ma et al., 2023c; Li et al., 2024d; Yang et al., 2024a; Zhao et al., 2024a; Li et al., 2024h; An et al., 2024; Ashkboos et al., 2024; Chen et al., 2023b; Xia et al., 2024; Guo et al., 2023b)] ] [Knowledge Distillation (§2.3.2) [White-Box KD (Zhang et al., 2023b; Gu et al., 2024a; Agarwal et al., 2024; Ko et al., 2024; Jha et al., 2024; Kim et al., 2024b; Padmanabhan et al., 2024); Black-Box KD (Peng et al., 2023b; Chen et al., 2023c; Wang et al., 2023f)] ] [Quantization (§2.3.3) [SqueezeLLM (Kim et al., 2023b); JSQ (Guo et al., 2024a); FrameQuant (Adepu et al., 2024); OneBit (Xu et al., 2024b); BiLLM (Huang et al., 2024b); LQER (Zhang et al., 2024c); I-LLM (Hu et al., 2024a); PV-Tuning (Malinovskii et al., 2024); BitNet (Wang et al., 2023e); BitNet b1.58 (Ma et al., 2024); PEQA (Kim et al., 2024a); QLoRA (Dettmers et al., 2024)] ] ] ] [Enhancement of SLMs (§3) [Training Methods for SLMs from Scratch (§3.1) [MobiLlama (Thawakar et al., 2024); MobileLLMs (Liu et al., 2024h); MindLLMs (Yang et al., 2023a); Tang et al. (2024a), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100 ] ] [Supervised Fine-Tuning (§3.2) [Alpaca (Taori et al., 2023); UltraChat (Ding et al., 2023); WizardLM (Xu et al., 2023b); SlimOrca (Lian et al., 2023); ShareGPT (Wang et al., 2024d); Capybara (Daniele and Suphavadeeprasit, 2023); Deita (Liu et al., 2023c); MetaMathQA (Yu et al., 2024a); MobileBERT (Sun et al., 2020); StableLM (Bellagente et al., 2024; Tow et al., [n. d.]); RLHF (Ouyang et al., 2022); DPO (Rafailov et al., 2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Data quality in Knowledge Distillation (§3.3) [TinyStory (Eldan and Li, 2023); Self-Amplify (Bhan et al., 2024); AS-ES Learning (Xi et al., 2024); Huang et al. ([n. d.]); Bhan et al. (2024); Tian et al. (2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Distillation techniques for enhancing SLM (§3.4) [GKD (Agarwal et al., 2024); DistiLLM (Ko et al., 2024); Adapt-and-Distill (Yao et al., 2021); Bit-level Inference Scaling Laws (Dettmers and Zettlemoyer, 2023), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Performance improvement through quantization (§3.5) [Bit-level Inference Scaling Laws (Dettmers and Zettlemoyer, 2023); BiLLM (Huang et al., 2024b); LLM.int8() (Dettmers et al., 2022); PB-LLM (Shang et al., 2023); OneBit (Xu et al., 2024b); BitNet (Wang et al., 2023e); LLM-QAT (Liu et al., 2023a), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Techniques in LLMs contributing SLMs (§3.6) [RAG for SLMs (Yazan et al., 2024; Liu et al., 2024g); MoE for SLMs (Kim et al., 2023a; Li et al., 2024e), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [Applications of SLMs (§4) [Task-specific SLM applications (§4.1) [SLM applications in QA (§4.1.1) [Phi-series (Gunasekar et al., 2023; Abdin et al., 2024); Orca 2 (Mitra et al., 2023); BioGPT-Large (Guo et al., 2023a); Stable Beluga 7B (Mahan et al., [n. d.]); Phogat et al. (2024); Hartill et al. (2023b); Gichamba et al. (2024); Jeong et al. (2023); Rationale Ranking (Hartill et al., 2023a)] ] [SLM applications in Coding (§4.1.2) [DeepSeek-Coder (Guo et al., 2024b); Phi-1 (Gunasekar et al., 2023); Phi-3.5-mini (Abdin et al., 2024); CodeGemma (Team, 2024a); CodeLlama (Roziere et al., 2023)] ] [SLM applications in Recommender Systems (§4.1.3) [PromptRec (Wu et al., 2024d); SLIM (Wang et al., 2024g); BiLLP (Shi et al., 2024); RecLoRA (Zhu et al., 2024); Wu et al. (2021)] ] [SLM applications in Web Search (§4.1.4) [Content encoder (Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020); H-ERNIE (Chu et al., 2022); Zou et al. (2022); Peng et al. (2023c); CoCondenser (Gao and Callan, 2022); Implicit Interaction ($I^{3}$) (Dong et al., 2023); InPars (Bonifacio et al., 2022); rewrite-retrieve-read (Ma et al., 2023d)] ] [SLM applications in Mobile-device (§4.1.5) [Octopus (Chen et al., 2024a); MobileAgent (Ding, 2024); Carreira et al. (2023); AutoDroid (Wen et al., 2024); Qin et al. (2023); Zhu et al. (2023b)] ] ] [Techniques during SLM Deployment (§4.2) [Memory efficiency (§4.2.1) [MoE-I2 (Yang et al., 2024f); MobileAIBench (Murthy et al., 2024); MobileLLM (Liu et al., 2024h); EdgeMoE (Yi et al., 2023); GEAR (Kang et al., 2024); HETLORA (Cho et al., 2024); MobiLlama (Thawakar et al., 2024)] ] [Computing efficiency (§4.2.2) [COST-EFF (Shen et al., 2022); MobileLLM (Liu et al., 2024h); MobiLlama (Thawakar et al., 2024); Merino (Zhao et al., 2024b); EdgeMoE (Yi et al., 2023); Nawrot et al. (2024); LLMCadLLM-Cad(Xu et al., 2023d); LinguaLinked (Zhao et al., 2023a)] ] ] ] [Models (§5) [Generic-domain SLMs (§5.1) [PhoneLM (Yi et al., 2024); Llama 3.2; Qwen (Bai et al., 2023; Yang et al., 2024h); Gemma (Team et al., 2024a, b); StableLM (Bellagente et al., 2024; Tow et al., [n. d.]); TinyLlama (Zhang et al., 2024j); OLMo (Groeneveld et al., 2024); H2O-Danube3 (Pfeiffer et al., 2024); Fox-1 (Team, 2024b); MiniCPM (Hu et al., 2024b); Phi (Gunasekar et al., 2023; Li et al., 2023b; Javaheripi et al., 2023; Abdin et al., 2024, 2024); BLOOM and BLOOMZ (Le Scao et al., 2023); Galactica (Taylor et al., 2022); OPT (Zhang et al., 2022); XGLM (Lin et al., 2022b); Megatron-gpt2 (Shoeybi et al., 2019); MINITRON (Muralidharan et al., 2024); LaMini-LM (Kim and Rush, 2016); FlanT5 (Chung et al., 2024) … , for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Specific-domain SLMs (§5.2) [Hippocrates (Acikgoz et al., 2024); BioMedLM (Bolton et al., 2024b); MentaLLaMA (Yang et al., 2024i); ChemLLM (Zhang et al., 2024e); SciGLM (Zhang et al., 2024d); AstroLLaMA (Nguyen et al., 2023); MindLLM (Yang et al., 2023a), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [SLMs for LLMs (§6) [SLM for reliable LLM generation (§6.1) [POLAR (Zhao et al., 2023b); SAPLMA (Azaria and Mitchell, 2023); SuperICL (Xu et al., 2023c); SuperContext (Yang et al., 2024j); Self-RAG (Asai et al., 2024); SKR (Wang et al., 2023c); SlimPLM (Tan et al., 2024); CRAG (Yan et al., 2024); Feng et al. (2023); Liu et al. (2024f); HaluAgent (Cheng et al., 2024); Toolformer (Schick et al., 2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for extracting LLM prompts (§6.2) [Prompt Stealing Attacks (Sha and Zhang, 2024); Output2prompt (Zhang et al., 2024f); Model Purifying (Li et al., 2024f); (Zhang et al., 2024b) , for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for fine-tuning LLMs (§6.3) [Weak-to-Strong Search (Zhou et al., 2024); Emulated Fine-tuning (Mitchell et al., 2024); CROSSLM (Deng et al., 2023); Swayamdipta et al. (2020); Mekala et al. (2024); Proxy-tuning (Liu et al., 2024d), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for LLM applications (§6.4) [SLCoLM (Tang et al., 2024b); HEF (Yang et al., 2024e); BLADE (Li et al., 2024a); Contrastive Decoding (Li et al., 2023e); “Train-Guide-Predict” (Tang et al., 2024b); Sennrich et al. (2023), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for LLM safety (§6.5) [Llama Guard (Inan et al., 2023); Kwon et al. (2024), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [SLM for LLM evaluation (§6.6) [SLIDE (Zhao et al., 2024d); Semantic uncertainty (Kuhn et al., 2023); Selfcheckgpt (Manakul et al., 2023); Proxylm (Anugraha et al., 2024); Factscore (Huang et al., 2023b), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [Synergy between SLMs and LLMs (§7) [Cloud-edge synergy (§7.1) [CoGenesis (Zhang et al., 2024h); Hao et al. (2024); Xu et al. (2024a); LLM-to-SLM (Bergner et al., 2024); Synergy of Thoughts (Shang et al., 2024); CROSSLM (Deng et al., 2023), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] [Task-centric synergy (§7.2) [$\alpha$-UMi (Shen et al., 2024a); SynCID (Liang et al., 2024); Filter-then-rerank Framework (Ma et al., 2023a); LLMCad (Xu et al., 2023d), for tree=text width=11cm, top color=customyellow!100, bottom color=customyellow!100] ] ] [Trustworthiness in SLMs (§8) [HELM (Liang et al., 2023); Do-Not-Answer (Wang et al., 2024f); PromptRobust (Zhu et al., 2023c); GLUE-X (Yang et al., 2023b); HaluEval (Li et al., 2023d); PrivLM-Bench (Li et al., 2024c); FFT (Cui et al., 2023); ROBBIE (Esiobu et al., 2023); TrustLLM (Sun et al., 2024b); RAmBLA (Bolton et al., 2024a); JailbreakBench (Chao et al., 2024); Xie et al. (2024b); OR-Bench (Cui et al., 2024); SORRY-Bench (Xie et al., 2024a); BeHonest (Chern et al., 2024); Hong et al. (2024); Nakka et al. (2024); RUPBench (Wang and Zhao, 2024), for tree=text width=15.53cm, top color=customyellow!100, bottom color=customyellow!100] ] [Future Directions (§9) ] ]
@@
The evolution of neural language models (LMs) from BERT’s (Devlin et al., 2019) pre-training and fine-tuning paradigm to T5’s (Raffel et al., 2020) pre-training plus prompting approach, and finally to GPT-3’s (Brown et al., 2020) pre-training plus in-context learning, has greatly enhanced natural language processing (NLP). These advancements have broadened NLP’s application across various fields, including language understanding (Wang et al., 2018), programming (Nam et al., 2024; Team, 2024a), recommendation systems (Wang et al., 2024g), information retrieval (Spatharioti et al., 2023; Chang et al., [n. d.]; Humeau et al., [n. d.]; Lu et al., 2020), mobile-device control (Ding, 2024), scientific discovery (Shojaee et al., 2024; Zhang et al., 2024e), medical question answering (Wang et al., 2023d; Bolton et al., 2024b), and legal question answering (Almeida et al., 2024). In particular, the recent emergence of proprietary commercial models including ChatGPT, Bard, and Claude, and open-sourced models such as Llama (Touvron et al., 2023a, b; Dubey et al., 2024) has led to rapid growth in the development of large language models (LLMs). Even though neural networks consistently improve on various tasks with longer training times, larger datasets, and increased model sizes—a phenomenon known as a neural scaling law (Kaplan et al., 2020), these models unpredictably exhibit a sudden acquisition of versatile abilities, termed ”emergent ability,” once they reach a critical scale threshold, thereby supporting the ”larger is better” trend. This ability is not present in small-scale models. For instance, the latest Llama-3.1 model with 405 billion parameters performs better in dialogue, logical reasoning, and programming compared to the smaller 7B counterpart (Dubey et al., 2024).
@@
Despite their prowess in complex tasks, LLMs’ huge parameters and computational needs impose significant limitations, hindering their adoption in many real-world applications. For example, the LLaMa 3.1 model with 405 billion parameters (Dubey et al., 2024), trained on 16K H100 GPUs for 54 days, requires about 202.5 GB of GPU memory using int4 precision and has large inference latency. These issues present several challenges in specific contexts: (1) LLMs are generally hosted in the cloud and used via cloud-based APIs due to the large GPU memory and computational cost. Users need to upload their data to query LLMs, raising data leakage and privacy concerns, especially in high-stake scenarios such as healthcare, finance, and e-commerce; (2) Driven by personal agents, on-device deployment is a critical requirement. Several factors, including cloud costs, latency, and privacy concerns, hinder the on-device processing of cloud-based LLMs, and direct deployment is impractical due to their high parameter and cache requirements, which often exceed the capabilities of devices such as mobile phones; (3) Their large parameter count can cause inference delays from seconds to minutes, unsuitable for real-time applications. For instance, Llama 2 7B takes approximately 84 seconds to process 100 tokens on benchmarks including HellaSwag, TruthfulQA, MMLU, and Arc_C when run on a smartphone equipped with a Snapdragon 685 processor (Thawakar et al., 2024); (4) To boost performance in specialized domains such as healthcare and law, where generic LLMs underperform, LLMs are often fine-tuned. However, this process is computationally expensive due to their large size. (5) Though general-purpose LLMs are powerful, many real-world applications require only specific abilities and domain knowledge, deploying general-purpose LLMs would be a waste of resources and such LLMs often cannot match the performance of models tailored for specific tasks (Phogat et al., 2024; Guo et al., 2023a; Jeong et al., 2023; Wang et al., 2024g; Chen et al., 2024a).
@@
Recently, small language models (SLMs) have shown great potential in alleviating these issues while achieving performance comparable to LLMs for domain-specific problems (Yang et al., 2024h; Team et al., 2024b; Bellagente et al., 2024; Pfeiffer et al., 2024; Zhang et al., 2024j; Groeneveld et al., 2024; Hu et al., 2024b; Abdin et al., 2024; Thawakar et al., 2024; Liu et al., 2024h). Owing to fewer parameters, SLMs excel in efficiency, cost, flexibility, and customization. They provide significant computational savings in pre-training and inference with reduced memory and storage needs, which is vital for applications requiring efficient resource use. These small models are especially effective in resource-limited settings, performing well on low-power devices such as edge devices. Besides, SLMs improve on-device processing by enhancing privacy, security, response times, and personalization. This supports advanced personal assistants and cloud-independent applications, boosting energy efficiency and reducing carbon emissions. For example, the Llama 3.2 models (1B & 3B) demonstrate that local processing enables immediate execution of prompts and responses (AI, 2024). This approach protects privacy by keeping sensitive data such as patient health information (PHI), business data, personal messages, and calendar details local, enhancing confidentiality. It also allows for precise control over which queries are processed on-device versus those requiring cloud-based models. Therefore, small language models are gaining increasing attention as alternatives to LLMs, as indicated in Figure 2, which shows that SLMs are downloaded more frequently than larger models in the Hugging Face community, and Figure 3, which illustrates the growing popularity of SLM releases over time.
@@
Typically, LMs that exhibit emergent abilities are classified as LLMs. However, the categorization of SLMs remains unclear. Studies vary in their contexts: some define SLMs as models with fewer than one billion parameters (Liu et al., 2024h), while others consider the term “small language model” relative to the larger counterparts (Wang et al., 2024g; Tang et al., 2024b; Lee et al., 2024b), with no consensus on a unified definition in the current landscape of LLMs. Research suggests SLMs for mobile devices, typically possessing around 6GB of memory, consist of sub-billion parameter models (Liu et al., 2024h), whereas others classify models with up to 10 billion parameters as small, noting their lack of emergent abilities (Fu et al., 2023). Given their use in resource-constrained environments and for specific tasks, we propose a generalized definition: Given specific tasks and resource constraints, we define SLMs as falling within a range where the lower bound is the minimum size at which the model exhibits emergent abilities for a specialized task, and the upper bound is the largest size manageable within limited resource conditions. This definition integrates various perspectives and addresses factors related to mobile computing and capability thresholds.
@@
Due to the growing demand for SLMs, extensive literature has emerged on various aspects of SLMs. For example, several resource-efficient techniques (Xu et al., 2024c) and training methods optimized for SLMs, such as quantization-aware training (Xu et al., 2024b; Wang et al., 2023e; Liu et al., 2023a) and selective architectural component choices (Qu et al., 2024a; Thawakar et al., 2024; Liu et al., 2024h), aim to enhance performance in specific applications (Phogat et al., 2024; Roziere et al., 2023; Wu et al., 2024d; Bonifacio et al., 2022; Chen et al., 2024a). These methods have led to the development of numerous open-source, general-purpose, and domain-specific SLMs (Yang et al., 2024h; Team et al., 2024b; Bellagente et al., 2024; Acikgoz et al., 2024; Bolton et al., 2024b; Zhang et al., 2024d). Beyond their inherent capabilities, SLMs can enhance LLMs by serving as modules or effective proxies (Zhao et al., 2023b; Xu et al., 2023a; Wu et al., 2024a; Sha and Zhang, 2024; Mitchell et al., 2024; Yang et al., 2024e). Furthermore, the complementary advantages of SLMs and LLMs can be leveraged collectively to better complete tasks (Zhang et al., 2024h; Deng et al., 2023; Shen et al., 2024a; Liang et al., 2024; Ma et al., 2023a). Despite the commendable performance of SLMs, it is crucial not to overlook their credibility issues, such as the risks of adversarial attacks, producing hallucinations, and privacy breaches (Dominguez-Olmedo et al., 2023; Hong et al., 2024; Egashira et al., 2024; Kumar et al., 2024; Nakka et al., 2024; Perez et al., 2023; Mo et al., 2024; Wang and Zhao, 2024; Kumar et al., 2024; Wang et al., 2023a; Yuan et al., 2024b; Liu et al., 2023b; Wang et al., 2021a, 2024b; Zhang et al., 2024i). However, currently, there is no comprehensive survey thoroughly exploring these works on SLMs in the era of LLMs. Therefore, this paper offers the first comprehensive survey that analyzes various aspects of SLMs in the LLM era and their future directions. The overview structure of our paper is shown in Figure 1. To summarize, our major contributions are:
@@
  • In Section 7, we survey two kinds of synergies between LLMs and SLMs: one involves cloud-based LLMs and local SLMs, while the other leverages the unique advantages of both to more effectively solve tasks.
@@
Concurrently with our survey, Lu et al. (2024) evaluate open-source SLMs, focusing on their architectures, datasets, algorithms, and on-device performance metrics such as inference latency and memory usage. Van Nguyen et al. (2024) delve into optimization strategies for SLMs, including model compression, pruning, and quantization. Chen and Varoquaux (2024) investigate how SLMs enhance LLMs and vice versa. In contrast, our survey offers a more comprehensive review with the following differences: (1) we present a detailed taxonomy of recent advancements in SLMs in the era of LLMs; (2) we define SLMs based on emergent capabilities and device specifications, which refines previous unclear definitions related to LLMs; (3) we discuss SLM applications, especially in on-device tasks and deployment, topics previously unexplored; (4) we examine domain-specific SLMs previously overlooked; and (5) we additionally consider the synergy between SLMs and LLMs.